# The Discovery of Little Red Dots in the Local Universe: Signatures of Cool Gas Envelopes

### v1

@@ 4.4 Variability
We examine the variability of all three objects in the optical () and infrared (WISE W1 and W2) bands. For the optical bands, we retrieve multi-epoch PSF-modeled photometry from the Pan-STARRS DR2 detection catalog. Over 1500 days (1300 days in the rest frame), we find no significant variability, with observed variation of 0.1–0.2 mag consistent with the scatter in non-variable sources (T. Simm et al., 2015). In the WISE W1 and W2 bands, we find variations of 0.25–0.5 mag over 5000 days (4000 days in the rest frame), also consistent with the typical scatter expected for non-variable sources. We thus conclude that no significant variability is detected in either the rest-frame optical or infrared bands, consistent with the lack of significant variability seen in high-redshift LRDs (M. Kokubo & Y. Harikane, 2024; Z. Zhang et al., 2025; X. Lin et al., 2025; W. L. Tee et al., 2025).
C. J. Burke et al. (2021) reported no significant variability in the broad H luminosity of J1025+1402 and J1047+0739 over a continuous 15-year period. The H luminosity range of J1025+1402 in C. J. Burke et al. (2021) is approximately – erg s-1, and for J1022+0841 it is – erg s-1, both well consistent with our measurements. They also presented V-band light curves from the Catalina Real-Time Transient Survey (CRTS; A. J. Drake et al., 2009) and -band light curves from the Zwicky Transient Facility (ZTF; F. J. Masci et al., 2019) for both objects. Although the limited quality of the light curves precluded definitive conclusions, they found potential low-level variability consistent with the stochastic damped random walk model of AGN variability (B. C. Kelly et al., 2009; C. L. MacLeod et al., 2010).
@@ 4.5 X-ray and Radio
C. Simmonds et al. (2016) reported net counts of Chandra for J1025+1402 in a net exposure time of 4937 s, placing an upper limit666Assuming a power law spectrum absorbed by Galactic . of . For J1047+0739, they reported net counts in 4782 s. We further examined the Chandra Source Catalog Release 2.1 (I. N. Evans et al., 2024), which includes more recent observations of J1047+0739, and found net counts in the 0.5–7 keV band over a 53 ks exposure. This corresponds to an X-ray luminosity of , although the detection is only significant at . For J1022+0841, no Chandra observations are currently available.
The comparison of the X-ray emission of J1025+1402 and J1047+0739 with that of typical type-1 AGNs is detailed in C. Simmonds et al. (2016). The upper limits or X-ray detections suggest that J1025+1402 and J1047+0739 are X-ray weak, exhibiting ratios at least 1–2 dex below those of typical type-1 AGNs with similar broad H luminosities (J. E. Greene & L. C. Ho, 2004; F. Panessa et al., 2006; L.-B. Desroches et al., 2009). This X-ray weakness is consistent with that seen in high-redshift LRDs (M. Yue et al., 2024; T. T. Ananna et al., 2024).
We examine the radio data from the VLA Sky Survey (M. Lacy et al., 2020), the TIFR GMRT Sky Survey (H. T. Intema et al., 2017), and the VLA FIRST Survey (R. H. Becker et al., 1995). We find no radio detections at 153 MHz, 1.4 GHz, and 3 GHz for any of the three sources. The corresponding upper limits for J1025+1402, J1047+0739, and J1022+0841 are: 13.8, 12.5, and 15.1 mJy at 153 MHz; 0.7, 0.7, and 0.8 mJy at 1.4 GHz; and 0.4, 0.5, and 0.4 mJy at 3 GHz, respectively. The upper limits of radio-loudness (, M. Sikora et al. 2007) for the three objects are estimated to be 8, 12, and 20, respectively. The current radio data are not deep enough to determine whether these sources are radio-quiet ().
@@ 4.6 Summary of Evidences as LRDs
As discussed above, the three objects at – satisfy all the defining criteria for LRDs. We summarize the similarity between the observed properties of the three local sources to high-redshift LRDs as follows:
@@
broad Balmer and Paschen emission lines, consistent with – BHs;
weak X-ray emission;
@@
high occurrence rate of Balmer and He I 10830 absorption.
Their broad H luminosities, erg s-1, fall within the range observed in high-redshift LRDs. If redshifted to , their expected apparent magnitudes in JWST/NIRCam F444W would be 25.7–26.8 mag. These values lie within the magnitude range observed for the diverse LRDs discovered at (e.g., J. E. Greene et al., 2024; D. D. Kocevski et al., 2024; H. B. Akins et al., 2024; D. D. Kocevski et al., 2025).
@@ 5 Emission Line Analysis
We utilize the rich series of Balmer and Paschen lines to measure the Balmer and Paschen decrements. For high-order hydrogen lines with insufficient S/N to detect reliable broad components, we fit them with a single Gaussian representing the narrow component, assuming that the broad components are buried in the noise. All measurements are based on the MODS spectra for consistency. As shown in Figure 7, the narrow Balmer lines from H to H12 in J1025+1402 and J1022+0841 closely match the intrinsic Case B recombination predictions (, ). The narrow Balmer lines in J1047+0739 are consistent with the SMC dust attenuation law (Y. C. Pei, 1992) with modest attenuation of . The exception is H and H8, which are blended with [Ne III] and He I 3889, respectively. The agreement indicates that the narrow hydrogen lines originate from regions with minimal dust attenuation. However, we caution that the decomposition of narrow and broad components is degenerate, and the strength of the potential underlying absorption remains uncertain.
@@
The broad H/H line ratios are in all three objects. Such large decrements may be attributed to either radiative transfer effects in high-density gas (e.g., collisional excitation and de-excitation, resonance scattering, etc.) or substantial dust attenuation. Notably, the measured decrements, both in the broad and narrow components, are well consistent with those observed in high-redshift LRDs (M. Brooks et al., 2025). These values are comparable to those expected for the SMC attenuation law with under Case B conditions. However, one caveat in attributing the broad-line decrements to dust is the survival of dust grains in the BLRs, which likely lies within the sublimation zone (e.g., P. G. Martin & G. J. Ferland, 1980; M. Kishimoto et al., 2007; C. M. Gaskell & P. Z. Harrington, 2018; A. Baskin & A. Laor, 2018). For comparison, type-1 AGNs in the SDSS typically show broad H/H ratios around 3.7, with the most extreme cases reaching 5 (D. Ilić et al., 2012). The much larger broad Balmer decrements observed in LRDs suggest that their BLRs may have extreme conditions with unusually dense gas or atypical dust properties, which differ substantially from those in typical type-1 AGNs (e.g., K. Inayoshi et al., 2022).
@@ 5.2 BPT Diagram and Metallicity
We estimate the BPT diagnostics (J. A. Baldwin et al., 1981) and the gas-phase metallicity of the narrow line regions (NLRs). To measure the fluxes of forbidden lines, we fit Gaussian profiles. Since the [O III] and lines are blended with [Fe II] emission, their fluxes are measured from spectra where the [Fe II] contribution has been subtracted using the best-fit models (see §5.3).
We show the positions of the three local LRDs on the [N II], [S II], and [O I] BPT diagrams in Figure 8. In the [N II] and [S II] diagrams, all three LRDs lie near the boundary between H II regions and AGN ionization. In the [O I] diagram, J1025+1402 and J1022+0841 fall within the Seyfert region, while J1047+0739 lies near the edge of the H II region. Their positions on the [N II] and [S II] diagrams also overlap with those of low-metallicity AGNs (L. J. Kewley et al., 2013; C. Simmonds et al., 2016). Notably, the locations of these local LRDs closely match that of the cosmic-noon LRD in I. Juodžbalis et al. (2024). This suggests that the narrow emission lines in both local and high-redshift LRDs are excited by similar mechanisms associated with AGN activity.
Additionally, we detect [Fe V] 4071 emission lines in all three objects. This high-ionization line further indicates the presence of a hard ionizing radiation field and AGN activity.
We determine the metallicities of the NLRs using the direct electron temperature () method. The of the O++ zone is derived from the ratio between [O III] and [O III], assuming cm-3. The estimated (O++) is K for J1025+1402, K for J1047+0739, and K for J1022+0841. The of the O+ zone is estimated using the empirical (O+)–(O++) relation from Y. I. Izotov et al. (2006). We apply a dust attenuation correction of to J1047+0739. As summarized in Table 1, all three objects exhibit metal-poor NLRs with 12 + ranging from 7.42 to 8.00. The values are broadly consistent with those reported by Y. I. Izotov & T. X. Thuan (2008) for J1025+1402 and J1047+0739.
We note that although the [O II] lines are detected in J1047+0739, they are significantly affected by recombination excitation enhanced by the hard ionizing radiation field (X. W. Liu et al., 2000; S. Tan et al., 2024). Therefore, we do not use them for (O+) diagnostics. The (O+) value directly derived from the [O II] and [O II] line ratios exceeds K, approximately twice that of (O++). This result further suggests an AGN origin for the narrow emission lines driven by hard ionization.
@@
5.3 [Fe II] Emission Analysis
The series of narrow [Fe II] emission lines observed in J1025+1402 and J1022+0841 stands in stark contrast to the broad Fe emission lines typically observed in type-1 AGNs (e.g., I Zw 1, M. P. Véron-Cetty et al., 2004). Though narrow optical [Fe II] lines have been reported in type-1 AGNs, they are blended with broad Fe lines and exhibit transitions and line ratios different from those in our sample; these narrow Fe lines are absent in type-2 AGNs (X.-B. Dong et al., 2010). To constrain the physical conditions such as temperature and density, we employ cloudy for photoionization modeling (C23, C. M. Gunasekera et al., 2023), with the latest Fe II atomic database (R. T. Smyth et al., 2019). We consider scenarios in which the gas is photoionized by AGN radiation. The modeling methodology is detailed in Appendix C.
The AGN photoionization models yield best-fit parameters for both J1025+1402 and J1022+0841 of , , and . The relative [Fe II] line ratios are not strongly sensitive to metallicity, so we adopt a fiducial value of without attempting to constrain it. The average temperature of the [Fe II]-emitting gas is approximately K, ranging from K in the inner region to K on the surface. As shown in Figure 9, the best-fit model reproduces the observed [Fe II] line ratios of J1025+1402 reasonably well, although some lines (e.g., [Fe II] 5278) remain imperfectly matched. A similar level of agreement is achieved for J1022+0841 using the same best-fit parameters (Figure C.1).
The inferred density lies between typical values for the BLRs and NLRs. The FWHMs of the [Fe II] lines are also intermediate ( km s-1 for J1025+1402 and km s-1 for J1022+0841), broader than the narrow H and [O III] (H –70 km s-1 in Binospec, [O III] not resolved in MODS). This suggests that the [Fe II] emission may arise from dense gas located outside the BLRs but inside the NLRs. We emphasize that our models are qualitative. Fe emission lines in AGNs are inherently complex and can be influenced by a variety of physical processes, making comprehensive theoretical modeling challenging (e.g., A. Sarkar et al., 2021; A. Pandey et al., 2025). Further refinement is required to fully understand the origin and nature of [Fe II] emission in the LRD population.
@@ 6 Case study: J1025+1402, “The Egg”
Hereafter, we nickname J1025+1402 as the Egg, inspired by its Ca-rich envelope, a temperature comparable to that of yellow supergiant stars, and its distinctive geometry. These features will be discussed in more detail below. An independent discovery and analysis of this object is also presented in Ji et al. (in prep).
@@ 6.1 Strong Metal Absorption: Evidence for a Cool Gas Envelope
The Egg exhibits exceptionally high-EW Na D 5890, 5896, K I 7665, 7699, Fe II 6148, 6244, 6456, and Ca triplets (8498, 8542, 8662; Ca T) absorption features in its optical continuum. We show their profiles in Figure 10 and summarize their properties in Table 2.
Absorber | EW | FWHM |
(Å) | (km s-1) | |
Na D doublets | ||
K I doublets | ||
Fe II triplets | ||
@@
6.1.1 Na D and K I Absorption
The Na D absorption in The Egg has an EW of Å and K I absorption shows an EW of Å. These two neutral alkali metals can be easily ionized by photons with low ionization potentials (5.14 eV for Na and 4.34 eV for K). This implies that the absorbing gas must be both cool and dense to preserve its neutral state. One possible origin for these absorbers is the cold, self-shielded intervening gas in the interstellar medium (ISM), with a minor contribution from late-type stars. However, the observed Na D EW in The Egg exceeds those typically found in such environments. The Na D EW in The Egg is higher than that observed in most galaxies and AGNs (typically Å; Y.-M. Chen et al., 2010; S. Cazzoli et al., 2016; D. S. N. Rupke et al., 2021; Y. Sun et al., 2024), and is only comparable to a few extreme IR-luminous LINERs (D. S. Rupke et al., 2005). In local galaxies and AGNs, high-EW Na D absorption typically traces gas outflows or inflows with velocities of several hundred km s-1 (S. Cazzoli et al., 2016; G. W. Roberts-Borsani & A. Saintonge, 2019; D. S. N. Rupke et al., 2021; Y. Sun et al., 2024), whereas in The Egg we do not observe significant velocity shifts.
Likewise, the K I EW is orders of magnitude larger than those found in the Galactic interstellar medium or in circumstellar gas clouds around young stellar objects, where values typically range from a few up to 200 mÅ (L. M. Hobbs, 1975; D. E. Welty & L. M. Hobbs, 2001; I. Pascucci et al., 2015). All of these lines of evidence point to a non-stellar and non-ISM origin.
6.1.2 Fe II Absorption
The optical Fe II absorption lines at 6148, 6244, and 6456 are rarely observed in galaxies or AGNs. These features have been reported in a Type Ia supernova (K. Maguire et al., 2023), but with large blueshifts of several thousand km s-1 and broad line widths. In contrast, the Fe II absorption seen in The Egg shows no significant velocity shift and narrow FWHMs of 400 km s-1. Although the origin of these Fe II absorption features remains uncertain, their co-occurrence with Na D, K I, and Ca T absorption may suggest a shared origin.
6.1.3 Ca T Absorption
The Ca T absorption at 8498, 8542 and 8662 Å in The Egg has an EW of Å. In Figure 11, we compare this value to typical Ca T EWs observed in stars, star-forming galaxies, and AGNs. The Ca T EW in The Egg exceeds those of most stars and all known galaxies and AGNs, where Ca T EWs are generally below 10 Å.
Only a few individual stars, such as yellow hypergiants or supergiants (S. V. Mallik, 1997; A. J. Cenarro et al., 2001), can exhibit comparable Ca T EWs. However, reproducing the observed Ca T EWs in the integrated galaxy light would require an unusually exotic stellar population composed solely of these stars.

### v2

@@ 4.4 Variability
We examine the variability of all three objects in the optical () and infrared (WISE W1 and W2) bands. For the optical bands, we retrieve multi-epoch PSF-modeled photometry from the Pan-STARRS DR2 detection catalog. Over 1500 days (1300 days in the rest frame), we find no significant variability, with observed variation of 0.1–0.2 mag consistent with systematic uncertainties as well as the scatter in non-variable sources (Simm2015). Following Lyu2019, we construct the WISE W1 (3.4 µm) and W2 (4.6 µm) band light curves from the WISE and NEOWISE (Mainzer2011) missions. In the W1 and W2 bands, we find variations of 0.25–0.5 mag over 5000 days (4000 days in the rest frame), also consistent with the typical scatter expected for non-variable sources. We thus conclude that no significant variability is detected in either the rest-frame optical or infrared bands, consistent with the lack of significant variability seen in high-redshift LRDs (Kokubo2024; ZZhang2025; Lin2025; Tee2025).
Burke2021 reported no significant variability in the broad H luminosity of J1025+1402 and J1047+0739 over a continuous 15-year period. The H luminosity range of J1025+1402 in Burke2021 is approximately – erg s-1, and for J1047+0739 it is – erg s-1, both well consistent with our measurements. They also presented V-band light curves from the Catalina Real-Time Transient Survey (CRTS; Drake2009) and -band light curves from the Zwicky Transient Facility (ZTF; Masci2019) for both objects. Although the limited quality of the light curves precluded definitive conclusions, they found potential low-level variability consistent with the stochastic damped random walk model of AGN variability (Kelly2009; MacLeod2010).
@@ 4.5 X-ray and Radio
Simmonds2016 reported net counts for J1025+1402 in the 0.5–7.0 keV band with Chandra, based on a net exposure time of 4937 s, placing an upper limit555Assuming a power law spectrum absorbed by Galactic . on the apparent X-ray luminosity of . For J1047+0739, they reported net counts in 4782 s. We further examined the Chandra Source Catalog Release 2.1 (Evans2024), which includes more recent observations of J1047+0739, and found net counts in the 0.5–7 keV band over a 53 ks exposure. This corresponds to an apparent X-ray luminosity of , although the detection is only significant at 666We use WebPIMMS to convert counts into fluxes. . For J1022+0841, no Chandra observations are currently available.
The comparison of the X-ray emission of J1025+1402 and J1047+0739 with that of typical type-1 AGNs is detailed in Simmonds2016. The upper limits or X-ray detections suggest that J1025+1402 and J1047+0739 are X-ray weak, exhibiting ratios at least 1–2 dex below those of typical type-1 AGNs with similar broad H luminosities (Greene2004; Panessa2006; Desroches2009). This X-ray weakness is consistent with that seen in high-redshift LRDs (Yue2024; Ananna2024).
We examine the radio data from the VLA Sky Survey (Lacy2020), the TIFR GMRT Sky Survey (Intema2017), and the VLA FIRST Survey (Becker1995). We find no radio detections at 153 MHz, 1.4 GHz, and 3 GHz for any of the three sources. The corresponding upper limits for J1025+1402, J1047+0739, and J1022+0841 are: 13.8, 12.5, and 15.1 mJy at 153 MHz; 0.7, 0.7, and 0.8 mJy at 1.4 GHz; and 0.4, 0.5, and 0.4 mJy at 3 GHz, respectively. The upper limits of radio-loudness (, Sikora2007) for the three objects are estimated to be 8, 12, and 20, respectively. The current radio data are not deep enough to determine whether these sources are radio-quiet ().
@@ 4.6 Summary of Evidences as LRDs
As discussed in this section and in the previous ones, the three objects at – satisfy all the defining criteria for LRDs. We summarize the similarity between the observed properties of the three local sources to high-redshift LRDs as follows:
@@
broad Balmer emission lines, consistent with – BHs;
weak X-ray emission; no radio detection in the current data;
@@
high occurrence rate of Balmer and He I 10833 absorption.
Their broad H luminosities, erg s-1, fall within the range observed in high-redshift LRDs (e.g., Matthee2024; Maiolino2024; Lin2024; Greene2024; Zhang2025; Hviding2025). If redshifted to , their expected apparent magnitudes in JWST/NIRCam F444W would be 25.7–26.8 mag. These values lie within the magnitude range observed for the diverse LRDs discovered at (e.g., Greene2024; Kocevski2024; Akins2024; Kocevski2025).
@@ 5 Emission Line Analysis
We utilize the rich series of Balmer and Paschen lines to measure the Balmer and Paschen decrements. For high-order hydrogen lines with insufficient S/N to detect reliable broad components, we fit them with a single Gaussian representing the narrow component, assuming that the broad components are buried in the noise. As shown in Figure 7, the narrow Balmer lines from H to H12 in J1025+1402 and J1022+0841 closely match the intrinsic Case B recombination predictions (, ; Brocklehurst1971; Hummer1987; Osterbrock2006). The narrow Balmer lines in J1047+0739 are consistent with the SMC dust attenuation law (Pei1992) with modest attenuation of . The exception is H and H8, which are blended with [Ne III] 3969 and He I 3890, respectively. The agreement indicates that the narrow hydrogen lines originate from regions with minimal dust attenuation. However, we caution that the decomposition of narrow and broad components is degenerate, and the strength of the potential underlying absorption remains uncertain.
@@
The broad H/H line ratios are for J1025+1402, for J1047+0739, and for J1022+0841. Such large decrements may be attributed to either radiative transfer effects in high-density gas (e.g., collisional excitation and de-excitation, resonance scattering, etc.) or substantial dust attenuation. Notably, the measured decrements of the broad Balmer lines are well consistent with those observed in high-redshift LRDs (Brooks2025; D'Eugenio2025_z7p04). These values are comparable to those expected for the SMC attenuation law with under Case B conditions. However, one caveat in attributing the broad-line decrements to dust is the survival of dust grains in the BLRs, which likely lies within the sublimation zone (e.g., Martin1980; Kishimoto2007; Gaskell2018; Baskin2018). For comparison, type-1 AGNs in the SDSS typically show broad H/H ratios around 3.7, with the most extreme cases reaching 5 (Ilic2012). The much larger broad Balmer decrements observed in LRDs suggest that their BLRs may have extreme conditions with unusually dense gas or atypical dust properties, which differ substantially from those in typical type-1 AGNs (e.g., Inayoshi2022).
@@ 5.2 BPT Diagram and Metallicity
We estimate the BPT diagnostics (Baldwin1981) and the gas-phase metallicity of the narrow line regions (NLRs). To measure the fluxes of forbidden lines, we fit Gaussian profiles. Since the [O III] and lines are blended with [Fe II] emission, their fluxes are measured from spectra where the [Fe II] contribution has been subtracted using the best-fit models (see §5.3).
We show the positions of the three local LRDs on the [N II], [S II], and [O I] BPT diagrams in Figure 8. In the [N II] and [S II] diagrams, all three LRDs lie near the boundary between H II regions and AGN ionization. In the [O I] diagram, J1025+1402 and J1022+0841 fall within the Seyfert region, while J1047+0739 lies near the edge of the H II region. Their positions on the [N II] and [S II] diagrams also overlap with those of low-metallicity AGNs (Kewley2013; Simmonds2016). Notably, the locations of these local LRDs closely match that of the cosmic-noon LRD in Juodzbalis2024. This suggests that the narrow emission lines in both local and high-redshift LRDs are excited by similar mechanisms associated with AGN activity.
We determine the metallicities of the NLRs using the direct electron temperature () method. The of the O++ zone is derived from the ratio between [O III] and [O III], assuming cm-3. The estimated (O++) is K for J1025+1402, K for J1047+0739, and K for J1022+0841. The of the O+ zone is estimated using the empirical (O+)–(O++) relation from Izotov2006. We apply a dust attenuation correction of to J1047+0739. As summarized in Table 1, all three objects exhibit metal-poor NLRs with 12 + ranging from 7.43 to 8.01. The values are broadly consistent with those reported by Izotov2008 for J1025+1402 and J1047+0739.
We note that although the [O II] lines are detected in J1047+0739, they are significantly affected by recombination excitation enhanced by the hard ionizing radiation field (Liu2000; Tan2024). Therefore, we do not use them for (O+) diagnostics. The (O+) value directly derived from the [O II] and [O II] line ratios exceeds K, more than twice the values of (O++) and the empirical (O+). This discrepancy suggests that [O II] is enhanced by recombination from the hard ionizing radiation field, and further supports an AGN origin for the narrow emission lines.
@@
5.3 [Fe II] (Fe II) Emission Analysis
We observed a series of narrow [Fe II] and Fe II emission lines in J1025+1402 and J1022+0841 (see Figures 9 and C.1), most of which are forbidden transitions. These lines stand in stark contrast to the broad Fe emission lines typically observed in type-1 AGNs (e.g., I Zw 1; VMP2004). Though narrow optical [Fe II] (Fe II) lines have been reported in type-1 AGNs, they are blended with broad Fe lines and exhibit transitions and line ratios different from those in our sample; these narrow Fe lines are absent in type-2 AGNs (Dong2010). To constrain the physical conditions such as temperature and density, we employ cloudy for photoionization modeling (C23, Gunasekera2023), with the latest Fe II atomic database (Smyth2019). We consider scenarios in which the gas is photoionized by AGN radiation. The modeling methodology is detailed in Appendix C.
The AGN photoionization models yield best-fit parameters for both J1025+1402 and J1022+0841 of , , and . The relative [Fe II] (Fe II) line ratios are not strongly sensitive to metallicity, so we adopt a fiducial value of without attempting to constrain it. The average temperature of the [Fe II]-emitting gas777For simplicity, we hereafter refer to the gas as [Fe II]-emitting, as most emission lines are forbidden [Fe II] transitions, although a few Fe II lines are also present. is approximately K, ranging from K in the inner region to K on the surface. As shown in Figure 9, the best-fit model reproduces the observed [Fe II] (Fe II) line ratios of J1025+1402 reasonably well, although some lines (e.g., Fe II 5277) remain imperfectly matched. A similar level of agreement is achieved for J1022+0841 using the same best-fit parameters (Figure C.1).
The inferred density lies between typical values for the BLRs and NLRs. The FWHMs of the [Fe II] (Fe II) lines are also intermediate ( km s-1 for J1025+1402 and km s-1 for J1022+0841), broader than the narrow H and [O III] (H km s-1 in Binospec, [O III] not resolved in MODS). This suggests that the [Fe II] (Fe II) emission may arise from dense gas located outside the BLRs but inside the NLRs. We emphasize that our models are qualitative. Fe emission lines in AGNs are inherently complex and can be influenced by a variety of physical processes, making comprehensive theoretical modeling challenging (e.g., Sarkar2021; Pandey2025). Further refinement is required to fully understand the origin and nature of [Fe II] (Fe II) emission in the LRD population.
@@ 6 Case study: J1025+1402, “The Egg”
Hereafter, we nickname J1025+1402 as the Egg, inspired by its Ca-rich envelope, a temperature comparable to that of yellow supergiant stars, and its distinctive geometry. These features will be discussed in more detail below. An independent discovery and analysis of this object is also presented in (Ji2025_lord).
@@ 6.1 Strong Metal Absorption: Evidence for a Cool Gas Envelope
6.1.1 Na D, K I, and Ca T Absorption
The Egg exhibits exceptionally high-EW Na D 5892, 5898, K I 7667, 7701, and Ca triplets (8500, 8544, 8665; Ca T) absorption features in its optical continuum. We show their profiles in Figure 10 and summarize their properties in Table 4.
| Absorber | EW | FWHM | |
| (Å) | (km s-1) | (km s-1) | |
| Na D doublets | |||
| K I doublets | |||
@@ 6.1.1 Na D, K I, and Ca T Absorption
The Na D absorption in The Egg has an EW of Å and K I absorption shows an EW of Å. These two neutral alkali metals can be easily ionized by photons with low ionization potentials (5.14 eV for Na and 4.34 eV for K). This implies that the absorbing gas must be both cool and dense to preserve its neutral state. One possible origin for these absorbers is the cold, self-shielded intervening gas in the interstellar medium (ISM), with a minor contribution from late-type stars. However, the observed Na D EW in The Egg exceeds those typically found in such environments. The Na D EW in The Egg is higher than that observed in most galaxies and AGNs (typically Å; Chen2010; Cazzoli2016; Rupke2021; Sun2024), and is only comparable to a few extreme IR-luminous LINERs (Rupke2005).
Likewise, the K I EW is orders of magnitude larger than those found in the Galactic interstellar medium or in circumstellar gas clouds around young stellar objects, where values typically range from a few up to 200 mÅ (Hobbs1975; Welty2001; Pascucci2015).
The Ca T absorption at 8500, 8544, and 8665 Å in The Egg has an EW of Å. In Figure 11, we compare this value to typical Ca T EWs observed in stars, star-forming galaxies, and AGNs. The Ca T EW in The Egg exceeds those of most stars and all known galaxies and AGNs, where Ca T EWs are generally below 10 Å.
Only a few individual stars, such as yellow hypergiants or supergiants (Mallik11997; Cenarro2001), can exhibit comparable Ca T EWs. However, reproducing the observed Ca T EWs in the integrated galaxy light would require an unusually exotic stellar population composed solely of these stars.
All the evidence above points to an origin that is neither stellar nor associated with the ISM. The distinct physical characteristics of these absorbers argue against a common origin within a single cloud. We discuss another possible origin in §6.1.3.
6.1.2 Other potential absorption lines
We identify additional weak absorption features in the optical spectrum of The Egg. We cross-matched the visual spectral troughs against three references: the stellar absorption-line list of Struve1934, the solar spectrum line list of Moore1966, and the theoretical line catalog from the vald stellar database (Piskunov1995; Ryabchikova2015). As shown in Figure 12, these spectral troughs likely correspond to low-ionization absorbers of Fe I, Ti II, Mg I, Na I, Ba II, Ca I, and Ti I. Some of the absorption features are also reported in Ji2025_lord. These species all have low ionization potentials: 7.9 eV for Fe I, 13.6 eV for Ti II, 7.6 eV for Mg I, 5.1 eV for Na I, 10.0 eV for Ba II, 6.1 eV for Ca I, and 6.8 eV for Ti I. These absorbers are commonly observed in cool G-to-K stars (e.g., Heiter2021). Additionally, CH G-band absorption is present, which was also highlighted in Ji2025_lord. This feature is characteristic of G–K type stars and traces cool, neutral regions that favor carbon-hydrogen molecular formation.